OpenStudy (anonymous):

Multiply -x – 2 by -x^2 + 4x – 2

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow (-x - 2)(-x^2 + 4x - 2) }\) \(\Large \color{purple}{\rightarrow -x(-x^2 + 4x - 2) - 2(-x^2 + 4x - 2) }\)

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x^3 - 4x^2 + 2x + 2x^2 - 8x + 4 }\)

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x^3 - 2x^2 - 6x + 4 }\)

OpenStudy (anonymous):

wasn't sure if i did it right but i do have the same answer. i did it a longer way. thanks

Parth (parthkohli):

Then you are correct, my friend.

OpenStudy (anonymous):

\[\begin{align} &&&-x-2\\ &\times&\color{green}{-x^2}&\color{red}{+4x}\color{blue}{-2}\\ \hline \\ &&&\color{blue}{+2x+4}\\ &&\color{red}{-4x^2}&\color{red}{-8x}\\ &+&\color{green}{x^3+2x^2}\\ \hline \\ &&x^3-2x^2&-6x+4 \end{align}\]

Parth (parthkohli):

@nbouscal How did you make that line?

OpenStudy (anonymous):

If you right-click my comment you should be able to select Show Math As TeX Commands and see everything I did

OpenStudy (anonymous):

his looks like the one in the books.

Parth (parthkohli):

@nbouscal I know but my browser doesn't support that.

OpenStudy (anonymous):

Oh, that sucks Parth, you should get Chrome. If you're asking about the horizontal line, it's just \hline, but I think you have to be within an {align} environment to use it.

Parth (parthkohli):

I have Chrome lol

OpenStudy (anonymous):

Oh, weird. It works on my Chrome.

Parth (parthkohli):

\hline doesn't work

OpenStudy (anonymous):

You have to have it between \begin {align} and \end {align} (no spaces)

OpenStudy (diyadiya):

\[ \begin{align} Like \ this \ =)\\ \hline \end{align} \]

Parth (parthkohli):

I couldn't make it wide though

OpenStudy (anonymous):

The line is as wide as the environment. You need to put in some filler to make the environment wider. \begin{align} \text{Add some filler text here.} \\ \hline \\ \end{align}

OpenStudy (anonymous):

BTW, Parth, there is a \(\LaTeX\) Practice study group on this site.

Parth (parthkohli):

\(\begin{align} \hline \text{I like it.} \end{align}\)

Parth (parthkohli):

I know, I'm a part of it.